# GEDAN: Learning the Edit Costs for Graph Edit Distance

### v1

@@ B.1 Results For Each Configuration
AIDS Morris et al. [2020] MUTAG Morris et al. [2020] PTC MR Morris et al. [2020] MODEL $\operatorname{RMSE}$ $\tau$ $\rho$ $\operatorname{RMSE}$ $\tau$ $\rho$ $\operatorname{RMSE}$ $\tau$ $\rho$ SimGNN (2019) Bai et al. [2019] $4.71\pm 0.36$ $\mathbf{0.563\pm 0.04}$ $\mathbf{0.793\pm 0.04}$ $14.20\pm 2.40$ $0.541\pm 0.07$ $0.798\pm 0.06$ $5.12\pm 0.40$ $\mathbf{0.698\pm 0.05}$ $\mathbf{0.881\pm 0.03}$ GREED (2022) Ranjan et al. [2022] $3.10\pm 0.20$ $0.349\pm 0.06$ $0.545\pm 0.06$ $5.66\pm 0.23$ $0.515\pm 0.07$ $0.705\pm 0.09$ $\mathbf{2.18\pm 0.08}$ $0.666\pm 0.06$ $0.849\pm 0.04$ ERIC (2022) Zhuo and Tan [2022] $6.30\pm 0.41$ $0.349\pm 0.06$ $0.545\pm 0.06$ $15.22\pm 2.78$ $0.512\pm 0.12$ $0.705\pm 0.09$ $5.83\pm 1.62$ $0.685\pm 0.05$ $0.871\pm 0.04$ GraphEdXXOR (2024) Jain et al. [2024] $4.26\pm 0.86$ $0.535\pm 0.07$ $0.730\pm 0.12$ $\mathbf{5.47\pm 0.19}$ $0.593\pm 0.02$ $\mathbf{0.861\pm 0.01}$ $3.38\pm 1.33$ $0.614\pm 0.16$ $0.784\pm 0.17$ U-GEDAN $3.54\pm 0.14$ $0.437\pm 0.06$ $0.652\pm 0.06$ $13.30\pm 0.85$ $\mathbf{0.611\pm 0.09}$ $0.852\pm 0.04$ $3.51\pm 0.31$ $0.649\pm 0.07$ $0.823\pm 0.05$ S-GEDAN $\mathbf{2.82\pm 0.19}$ $0.484\pm 0.05$ $0.694\pm 0.05$ $6.01\pm 0.66$ $0.608\pm 0.09$ $0.855\pm 0.04$ $3.18\pm 0.65$ $0.655\pm 0.08$ $0.826\pm 0.07$
@@
AIDS Morris et al. [2020] MUTAG Morris et al. [2020] PTC MR Morris et al. [2020] MODEL $\operatorname{RMSE}$ $\tau$ $\rho$ $\operatorname{RMSE}$ $\tau$ $\rho$ $\operatorname{RMSE}$ $\tau$ $\rho$ SimGNN (2019) Bai et al. [2019] $6.00\pm 0.45$ $\mathbf{0.583\pm 0.04}$ $\mathbf{0.803\pm 0.03}$ $6.53\pm 1.19$ $\mathbf{0.615\pm 0.05}$ $0.827\pm 0.05$ $4.93\pm 0.60$ $\mathbf{0.775\pm 0.05}$ $\mathbf{0.928\pm 0.03}$ GREED (2022) Ranjan et al. [2022] $\mathbf{2.92\pm 0.09}$ $0.334\pm 0.07$ $0.571\pm 0.07$ $4.48\pm 0.86$ $0.585\pm 0.04$ $0.676\pm 0.06$ $\mathbf{2.06\pm 0.18}$ $0.711\pm 0.05$ $0.881\pm 0.05$ ERIC (2022) Zhuo and Tan [2022] $6.16\pm 0.48$ $0.575\pm 0.04$ $0.797\pm 0.03$ $6.02\pm 0.30$ $0.591\pm 0.06$ $\mathbf{0.849\pm 0.02}$ $6.34\pm 1.29$ $0.762\pm 0.03$ $0.902\pm 0.03$ GraphEdXXOR (2024) Jain et al. [2024] $7.14\pm 0.14$ $0.537\pm 1.36$ $0.739\pm 0.06$ $4.08\pm 0.82$ $0.569\pm 0.11$ $0.725\pm 0.08$ $3.44\pm 1.79$ $0.671\pm 0.19$ $0.820\pm 0.18$ U-GEDAN $3.91\pm 0.09$ $0.482\pm 0.06$ $0.698\pm 0.05$ $6.34\pm 0.80$ $0.595\pm 0.11$ $0.813\pm 0.05$ $3.44\pm 0.24$ $0.742\pm 0.05$ $0.897\pm 0.03$ S-GEDAN $3.52\pm 0.14$ $0.511\pm 0.04$ $0.721\pm 0.04$ $\mathbf{3.36\pm 0.49}$ $0.587\pm 0.09$ $0.829\pm 0.04$ $2.85\pm 0.47$ $0.770\pm 0.02$ $0.917\pm 0.02$
@@
AIDS Morris et al. [2020] MUTAG Morris et al. [2020] PTC MR Morris et al. [2020] MODEL $\operatorname{RMSE}$ $\tau$ $\rho$ $\operatorname{RMSE}$ $\tau$ $\rho$ $\operatorname{RMSE}$ $\tau$ $\rho$ SimGNN (2019) Bai et al. [2019] $13.20\pm 3.83$ $0.497\pm 0.04$ $0.670\pm 0.06$ $10.72\pm 0.28$ $0.469\pm 0.04$ $0.618\pm 0.04$ $12.91\pm 1.24$ $0.579\pm 0.07$ $0.760\pm 0.07$ GREED (2022) Ranjan et al. [2022] $\mathbf{3.05\pm 0.16}$ $0.467\pm 0.05$ $0.663\pm 0.06$ $4.96\pm 0.78$ $\mathbf{0.592\pm 0.05}$ $0.659\pm 0.06$ $\mathbf{2.85\pm 0.34}$ $\mathbf{0.676\pm 0.07}$ $\mathbf{0.847\pm 0.06}$ ERIC (2022) Zhuo and Tan [2022] $12.24\pm 0.83$ $0.467\pm 0.05$ $0.662\pm 0.05$ $8.12\pm 0.59$ $0.573\pm 0.07$ $0.809\pm 0.03$ $13.13\pm 2.02$ $0.611\pm 0.05$ $0.787\pm 0.05$ GraphEdXXOR (2024) Jain et al. [2024] $10.76\pm 0.50$ $\mathbf{0.586\pm 0.01}$ $\mathbf{0.803\pm 0.02}$ $\mathbf{4.07\pm 0.90}$ $0.551\pm 0.09$ $0.715\pm 0.11$ $5.39\pm 0.27$ $0.637\pm 0.02$ $0.815\pm 0.02$ U-GEDAN $6.72\pm 0.18$ $0.398\pm 0.05$ $0.608\pm 0.04$ $15.12\pm 1.18$ $0.569\pm 0.08$ $\mathbf{0.830\pm 0.03}$ $6.29\pm 0.40$ $0.639\pm 0.06$ $0.815\pm 0.05$ S-GEDAN $5.88\pm 0.28$ $0.426\pm 0.05$ $0.641\pm 0.04$ $4.08\pm 0.56$ $0.582\pm 0.08$ $0.825\pm 0.04$ $5.42\pm 1.03$ $0.664\pm 0.07$ $0.835\pm 0.06$
@@
AIDS Morris et al. [2020] MUTAG Morris et al. [2020] PTC MR Morris et al. [2020] MODEL $\operatorname{RMSE}$ $\tau$ $\rho$ $\operatorname{RMSE}$ $\tau$ $\rho$ $\operatorname{RMSE}$ $\tau$ $\rho$ SimGNN (2019) Bai et al. [2019] $9.55\pm 0.80$ $0.542\pm 0.05$ $0.770\pm 0.05$ $8.87\pm 0.81$ $\mathbf{0.641\pm 0.02}$ $\mathbf{0.848\pm 0.02}$ $10.57\pm.98$ $0.678\pm 0.04$ $0.850\pm 0.03$ GREED (2022) Ranjan et al. [2022] $\mathbf{3.06\pm 0.22}$ $0.484\pm 0.07$ $0.698\pm 0.06$ $4.77\pm 0.67$ $0.596\pm 0.04$ $0.648\pm 0.04$ $\mathbf{3.25\pm 0.19}$ $\mathbf{0.696\pm 0.05}$ $\mathbf{0.858\pm 0.03}$ ERIC (2022) Zhuo and Tan [2022] $10.90\pm 0.69$ $0.507\pm 0.05$ $0.725\pm 0.05$ $9.29\pm 0.60$ $0.413\pm 0.07$ $0.596\pm 0.05$ $11.02\pm 1.47$ $0.613\pm 0.08$ $0.779\pm 0.09$ GraphEdXXOR (2024) Jain et al. [2024] $8.56\pm 0.28$ $\mathbf{0.596\pm 0.03}$ $\mathbf{0.827\pm 0.01}$ $\mathbf{4.50\pm 0.61}$ $0.610\pm 0.04$ $0.816\pm 0.05$ $5.16\pm 0.48$ $0.664\pm 0.04$ $0.844\pm 0.04$ U-GEDAN $5.77\pm 0.12$ $0.432\pm 0.04$ $0.675\pm 0.04$ $7.50\pm 0.72$ $0.604\pm 0.10$ $0.827\pm 0.06$ $5.75\pm 0.52$ $0.638\pm 0.06$ $0.831\pm 0.04$ S-GEDAN $5.17\pm 0.26$ $0.479\pm 0.04$ $0.718\pm 0.04$ $4.82\pm 0.58$ $0.599\pm 0.10$ $0.825\pm 0.06$ $5.27\pm 1.04$ $0.681\pm 0.08$ $0.856\pm 0.06$
@@
AIDS Morris et al. [2020] MUTAG Morris et al. [2020] PTC MR Morris et al. [2020] MODEL $\operatorname{RMSE}$ $\tau$ $\rho$ $\operatorname{RMSE}$ $\tau$ $\rho$ $\operatorname{RMSE}$ $\tau$ $\rho$ SimGNN (2019) Bai et al. [2019] $10.14\pm 0.72$ $\mathbf{0.562\pm 0.04}$ $0.763\pm 0.04$ $7.96\pm 1.38$ $\mathbf{0.642\pm 0.05}$ $\mathbf{0.840\pm 0.05}$ $9.32\pm 0.68$ $0.717\pm 0.05$ $\mathbf{0.890\pm 0.03}$ GREED (2022) Ranjan et al. [2022] $\mathbf{3.14\pm 0.22}$ $0.509\pm 0.06$ $0.690\pm 0.07$ $\mathbf{4.51\pm 0.40}$ $0.630\pm 0.03$ $0.697\pm 0.04$ $\mathbf{2.99\pm 0.46}$ $\mathbf{0.719\pm 0.07}$ $0.872\pm 0.05$ ERIC (2022) Zhuo and Tan [2022] $11.80\pm 0.33$ $0.461\pm 0.04$ $0.622\pm 0.04$ $9.76\pm 1.27$ $0.560\pm 0.05$ $0.779\pm 0.02$ $11.69\pm 1.37$ $0.652\pm 0.08$ $0.769\pm 0.09$ GraphEdXXOR (2024) Jain et al. [2024] $9.56\pm 0.25$ $0.545\pm 0.01$ $\mathbf{0.832\pm 0.02}$ $4.92\pm 0.85$ $0.529\pm 0.06$ $0.688\pm 0.07$ $4.64\pm 0.25$ $0.712\pm 0.02$ $0.879\pm 0.01$ U-GEDAN $6.00\pm 0.10$ $0.433\pm 0.04$ $0.658\pm 0.04$ $8.65\pm 0.94$ $0.513\pm 0.05$ $0.741\pm 0.02$ $5.56\pm 0.55$ $0.650\pm 0.06$ $0.839\pm 0.05$ S-GEDAN $5.50\pm 0.25$ $0.486\pm 0.04$ $0.705\pm 0.04$ $4.91\pm 0.47$ $0.521\pm 0.06$ $0.755\pm 0.03$ $5.51\pm 1.10$ $0.693\pm 0.06$ $0.848\pm 0.07$
Appendix C Details of Edit Cost Optimization

### v2

@@ D.2 Results For Each Configuration
AIDS MUTAG PTC MR MODEL $\operatorname{RMSE}~(\downarrow)$ $\tau~(\uparrow)$ $\rho~(\uparrow)$ $\operatorname{RMSE}~(\downarrow)$ $\tau~(\uparrow)$ $\rho~(\uparrow)$ $\operatorname{RMSE}~(\downarrow)$ $\tau~(\uparrow)$ $\rho~(\uparrow)$ S SimGNN [1] $3.47\pm 0.33$ $\mathbf{0.571\pm 0.05}$ $\mathbf{0.764\pm 0.04}$ $7.96\pm 1.04$ $0.562\pm 0.06$ $0.802\pm 0.03$ $4.08\pm 0.37$ $\underline{0.708\pm 0.04}$ $\underline{0.884\pm 0.03}$ EGSC [44] $\underline{2.94\pm 0.12}$ $0.463\pm 0.02$ $0.638\pm 0.03$ $5.56\pm 0.38$} $0.553\pm 0.03$ $0.754\pm 0.04$ $\underline{2.37\pm 0.21}$ $\mathbf{0.729\pm 0.02}$ $\mathbf{0.899\pm 0.03}$ GREED [45] $3.10\pm 0.20$ $0.349\pm 0.06$ $0.545\pm 0.06$ $5.66\pm 0.23$ $0.515\pm 0.07$ $0.705\pm 0.09$ $\mathbf{2.18\pm 0.08}$ $0.666\pm 0.06$ $0.849\pm 0.04$ ERIC [69] $3.66\pm 0.34$ $0.425\pm 0.05$ $0.656\pm 0.03$ $8.65\pm 1.28$ $0.525\pm 0.06$ $0.775\pm 0.06$ $4.32\pm 1.10$ $0.706\pm 0.04$ $0.865\pm 0.04$ GraphEdXXOR [28] $3.58\pm 0.46$ $\underline{0.557\pm 0.04}$ $\underline{0.745\pm 0.08}$ $\underline{5.33\pm 0.14}$ $\underline{0.605\pm 0.02}$ $\underline{0.868\pm 0.01}$ $3.30\pm 0.53$ $0.651\pm 0.06$ $0.835\pm 0.06$ GEDAN($\lambda=0$) $\mathbf{2.82\pm 0.19}$ $0.484\pm 0.05$ $0.694\pm 0.05$ $\mathbf{5.20\pm 0.15}$ $\mathbf{0.650\pm 0.04}$ $\mathbf{0.887\pm 0.01}$ $3.18\pm 0.65$ $0.655\pm 0.08$ $0.826\pm 0.07$ U Approximated GED [20] $\underline{3.89}$ $0.372$ $0.542$ $\underline{16.46}$ $\underline{0.560}$ $\underline{0.757}$ $\mathbf{3.49}$ $\underline{0.630}$ $\underline{0.797}$ EUGENE [6] $4.32\pm 0.16$ $\mathbf{0.505\pm 0.03}$ $\mathbf{0.658\pm 0.05}$ $18.68\pm 0.80$ $0.507\pm 0.09$ $0.684\pm 0.07$ $5.15\pm 0.76$ $0.625\pm 0.06$ $0.763\pm 0.05$ GEDAN($\lambda=1$) $\mathbf{3.54\pm 0.14}$ $\underline{0.437\pm 0.06}$ $\underline{0.652\pm 0.06}$ $\mathbf{13.30\pm 0.85}$ $\mathbf{0.611\pm 0.09}$ $\mathbf{0.852\pm 0.04}$ $\underline{3.51\pm 0.31}$ $\mathbf{0.649\pm 0.07}$ $\mathbf{0.823\pm 0.05}$
@@
AIDS MUTAG PTC MR MODEL $\operatorname{RMSE}~(\downarrow)$ $\tau~(\uparrow)$ $\rho~(\uparrow)$ $\operatorname{RMSE}~(\downarrow)$ $\tau~(\uparrow)$ $\rho~(\uparrow)$ $\operatorname{RMSE}~(\downarrow)$ $\tau~(\uparrow)$ $\rho~(\uparrow)$ S SimGNN [1] $5.82\pm 0.36$ $\mathbf{0.591\pm 0.03}$ $\mathbf{0.810\pm 0.02}$ $6.02\pm 0.76$ $\mathbf{0.631\pm 0.04}$ $0.821\pm 0.04$ $4.08\pm 0.51$ $0.755\pm 0.04$ $0.915\pm 0.03$} EGSC [44] $\underline{3.36\pm 0.11}$ $0.512\pm 0.02$ $0.696\pm 0.02$ $\mathbf{2.47\pm 0.21}$ $\underline{0.604\pm 0.03}$ $0.808\pm 0.03$ $\underline{2.52\pm 0.19}$ $\mathbf{0.782\pm 0.02}$ $\mathbf{0.931\pm 0.02}$ GREED [45] $\mathbf{2.92\pm 0.09}$ $0.334\pm 0.07$ $0.571\pm 0.07$ $4.48\pm 0.86$ $0.585\pm 0.04$ $0.676\pm 0.06$ $\mathbf{2.06\pm 0.18}$ $0.711\pm 0.05$ $0.881\pm 0.05$ ERIC [69] $5.25\pm 0.42$ $0.545\pm 0.03$} $\underline{0.763\pm 0.05}$ $5.78\pm 0.35$ $0.588\pm 0.07$ $\mathbf{0.831\pm 0.03}$ $4.94\pm 0.89$ $0.763\pm 0.03$ $0.911\pm 0.02$ GraphEdXXOR [28] $4.87\pm 0.10$ $\underline{0.557\pm 0.06}$ $0.745\pm 0.05$ $3.85\pm 0.51$ $0.574\pm 0.08$ $0.759\pm 0.07$ $3.12\pm 0.22$ $0.694\pm 0.07$ $0.833\pm 0.06$ GEDAN($\lambda=0$) $3.52\pm 0.14$ $0.511\pm 0.04$ $0.721\pm 0.04$ $\underline{3.36\pm 0.49}$ $0.587\pm 0.09$ $\underline{0.829\pm 0.04}$ $2.85\pm 0.47$ $\underline{0.770\pm 0.02}$ $\underline{0.917\pm 0.02}$ U Approximated GED [20] 6.29{6.29} 0.366{0.366} 0.532{0.532} $\underline{8.01}$ $\underline{0.442}$ $\underline{0.589}$ $\underline{4.42}$ 0.661{0.661} 0.829{0.829} EUGENE [6] $\underline{5.39\pm 0.19}$ $\mathbf{0.514\pm 0.02}$ $\underline{0.691\pm 0.04}$ $7.02\pm 1.37$ $0.402\pm 0.18$ $0.588\pm 0.13$ $5.14\pm 0.54$ $\underline{0.709\pm 0.03}$ $\underline{0.848\pm 0.02}$ GEDAN($\lambda=1$) $\mathbf{3.91\pm 0.09}$ $\underline{0.482\pm 0.06}$ $\mathbf{0.698\pm 0.05}$ $\mathbf{6.34\pm 0.80}$ $\mathbf{0.595\pm 0.11}$ $\mathbf{0.813\pm 0.05}$ $\mathbf{3.44\pm 0.24}$ $\mathbf{0.742\pm 0.05}$ $\mathbf{0.897\pm 0.03}$
@@
AIDS MUTAG PTC MR MODEL $\operatorname{RMSE}~(\downarrow)$ $\tau~(\uparrow)$ $\rho~(\uparrow)$ $\operatorname{RMSE}~(\downarrow)$ $\tau~(\uparrow)$ $\rho~(\uparrow)$ $\operatorname{RMSE}~(\downarrow)$ $\tau~(\uparrow)$ $\rho~(\uparrow)$ S SimGNN [1] $7.32\pm 0.91$ $\underline{0.523\pm 0.05}$ $\underline{0.706\pm 0.04}$ $6.55\pm 0.23$ $0.513\pm 0.05$ $0.722\pm 0.03$ $6.84\pm 1.05$ $0.605\pm 0.06$ $0.767\pm 0.04$ EGSC [44] $6.15\pm 0.16$ $0.413\pm 0.01$ $0.579\pm 0.01$ $\mathbf{2.96\pm 0.13}$ $0.572\pm 0.04$ $0.784\pm 0.03$ $\underline{5.03\pm 0.21}$ $\underline{0.681\pm 0.02}$ $\mathbf{0.860\pm 0.02}$ GREED [45] $\mathbf{3.05\pm 0.16}$ $0.467\pm 0.05$ $0.663\pm 0.06$ $4.96\pm 0.78$ $\mathbf{0.592\pm 0.05}$ $0.659\pm 0.06$ $\mathbf{2.85\pm 0.34}$ $0.676\pm 0.07$} $0.847\pm 0.06$} ERIC [69] $6.74\pm 0.53$ $0.438\pm 0.04$ $0.631\pm 0.05$ $5.70\pm 0.31$ $0.560\pm 0.05$ $\underline{0.793\pm 0.04}$ $6.11\pm 1.12$ $0.630\pm 0.05$ $0.795\pm 0.04$ GraphEdXXOR [28] $\underline{5.45\pm 0.43}$ $\mathbf{0.591\pm 0.02}$ $\mathbf{0.811\pm 0.02}$ $\underline{3.98\pm 0.65}$ $0.571\pm 0.08$ $0.753\pm 0.09$ $5.10\pm 0.21$ $\mathbf{0.695\pm 0.02}$ $\underline{0.848\pm 0.03}$ GEDAN($\lambda=0$) $5.88\pm 0.28$ $0.426\pm 0.05$ $0.641\pm 0.04$ $4.08\pm 0.56$ $\underline{0.582\pm 0.08}$ $\mathbf{0.825\pm 0.04}$ $5.42\pm 1.03$ $0.664\pm 0.07$ $0.835\pm 0.06$ U Approximated GED [20] $\underline{8.15}$ 0.328{0.328} 0.495{0.495} $\mathbf{8.01}$ $\underline{0.444}$ $\underline{0.613}$ $\underline{6.30}$ $\mathbf{0.643}$ $\underline{0.804}$ EUGENE [6] $12.50\pm 0.94$ $\mathbf{0.437\pm 0.04}$ $\underline{0.564\pm 0.05}$ $\underline{9.18\pm 1.56}$ $0.372\pm 0.20$ $0.553\pm 0.15$ $10.85\pm 0.99$ $0.628\pm 0.07$ $0.764\pm 0.05$ GEDAN($\lambda=1$) $\mathbf{6.72\pm 0.18}$ $\underline{0.398\pm 0.05}$ $\mathbf{0.608\pm 0.04}$ $15.12\pm 1.18$ $\mathbf{0.569\pm 0.08}$ $\mathbf{0.830\pm 0.03}$ $\mathbf{6.29\pm 0.40}$ $\underline{0.639\pm 0.06}$ $\mathbf{0.815\pm 0.05}$
@@
AIDS MUTAG PTC MR MODEL $\operatorname{RMSE}~(\downarrow)$ $\tau~(\uparrow)$ $\rho~(\uparrow)$ $\operatorname{RMSE}~(\downarrow)$ $\tau~(\uparrow)$ $\rho~(\uparrow)$ $\operatorname{RMSE}~(\downarrow)$ $\tau~(\uparrow)$ $\rho~(\uparrow)$ S SimGNN [1] $7.05\pm 0.72$ $\underline{0.525\pm 0.04}$ $\underline{0.755\pm 0.04}$ $6.31\pm 0.66$ $\underline{0.641\pm 0.02}$ $0.833\pm 0.03$} $7.11\pm 0.58$ $0.658\pm 0.04$ $0.825\pm 0.06$ EGSC [44] $5.21\pm 0.17$ $0.474\pm 0.03$ $0.651\pm 0.03$ $\mathbf{3.27\pm 0.16}$ $\mathbf{0.672\pm 0.02}$ $\mathbf{0.863\pm 0.02}$ $\underline{4.89\pm 0.62}$ $\mathbf{0.709\pm 0.02}$ $\mathbf{0.880\pm 0.02}$ GREED [45] $\mathbf{3.06\pm 0.22}$ $0.484\pm 0.07$ $0.698\pm 0.06$ $4.77\pm 0.67$ $0.596\pm 0.04$ $0.648\pm 0.04$ $\mathbf{3.25\pm 0.19}$ $\underline{0.696\pm 0.05}$ $0.858\pm 0.03$} ERIC [69] $6.17\pm 0.55$ $0.507\pm 0.04$ $0.728\pm 0.05$ $5.54\pm 0.65$ $0.581\pm 0.06$ $0.652\pm 0.06$ $5.81\pm 1.10$ $0.652\pm 0.08$ $0.815\pm 0.04$ GraphEdXXOR [28] $\underline{4.65\pm 0.32}$ $\mathbf{0.605\pm 0.03}$ $\mathbf{0.835\pm 0.02}$ $\underline{4.44\pm 0.65}$ $0.624\pm 0.04$ $\underline{0.838\pm 0.05}$ $5.25\pm 0.68$ $0.666\pm 0.05$ $\underline{0.861\pm 0.05}$ GEDAN($\lambda=0$) $5.17\pm 0.26$} $0.479\pm 0.04$ $0.718\pm 0.04$ $4.82\pm 0.58$ $0.599\pm 0.10$ $0.825\pm 0.06$ $5.27\pm 1.04$ $0.681\pm 0.08$ $0.856\pm 0.06$ U Approximated GED [20] $\underline{8.22}$ 0.181{0.181} 0.269{0.269} $\underline{9.24}$ 0.278{0.278} 0.416{0.416} 10.96{10.96} 0.290{0.290} 0.350{0.350} EUGENE [6] $9.50\pm 0.56$ $\mathbf{0.449\pm 0.03}$ $\underline{0.659\pm 0.04}$ $10.96\pm 1.43$ $\underline{0.467\pm 0.13}$ $\underline{0.671\pm 0.11}$ $\underline{9.54\pm 1.26}$ $\underline{0.622\pm 0.09}$ $\underline{0.796\pm 0.07}$ GEDAN($\lambda=1$) $\mathbf{5.77\pm 0.12}$ $\underline{0.432\pm 0.04}$ $\mathbf{0.675\pm 0.04}$ $\mathbf{7.50\pm 0.72}$ $\mathbf{0.604\pm 0.10}$ $\mathbf{0.827\pm 0.06}$ $\mathbf{5.75\pm 0.52}$ $\mathbf{0.638\pm 0.06}$ $\mathbf{0.831\pm 0.04}$
@@
AIDS MUTAG PTC MR MODEL $\operatorname{RMSE}~(\downarrow)$ $\tau~(\uparrow)$ $\rho~(\uparrow)$ $\operatorname{RMSE}~(\downarrow)$ $\tau~(\uparrow)$ $\rho~(\uparrow)$ $\operatorname{RMSE}~(\downarrow)$ $\tau~(\uparrow)$ $\rho~(\uparrow)$ S SimGNN [1] $5.98\pm 0.63$ $\underline{0.571\pm 0.06}$ $\underline{0.788\pm 0.05}$ $4.77\pm 0.85$ $\underline{0.656\pm 0.04}$ $0.837\pm 0.03$} $4.64\pm 0.55$ $0.718\pm 0.03$ $\underline{0.885\pm 0.03}$ EGSC [44] $5.47\pm 0.49$} $0.479\pm 0.02$ $0.653\pm 0.03$ $\mathbf{2.76\pm 0.16}$ $\mathbf{0.667\pm 0.02}$ $\mathbf{0.864\pm 0.02}$ $\underline{4.50\pm 0.35}$ $\mathbf{0.731\pm 0.04}$ $\mathbf{0.899\pm 0.03}$ GREED [45] $\mathbf{3.14\pm 0.22}$ $0.509\pm 0.06$ $0.690\pm 0.07$ $4.51\pm 0.40$} $0.630\pm 0.03$ $0.697\pm 0.04$ $\mathbf{2.99\pm 0.46}$ $\underline{0.719\pm 0.07}$ $0.872\pm 0.05$ ERIC [69] $5.36\pm 0.43$ $0.461\pm 0.04$ $0.666\pm 0.03$ $4.56\pm 0.77$ $0.582\pm 0.03$ $0.805\pm 0.03$ $5.05\pm 0.83$ $0.687\pm 0.04$ $0.865\pm 0.05$ GraphEdXXOR [28] $\underline{4.65\pm 0.48}$ $\mathbf{0.605\pm 0.02}$ $\mathbf{0.835\pm 0.02}$ $\underline{4.44\pm 0.76}$ $0.624\pm 0.07$ $\underline{0.838\pm 0.03}$ $5.25\pm 0.32$ $0.666\pm 0.03$ $0.861\pm 0.04$ GEDAN($\lambda=0$) $5.50\pm 0.25$ $0.486\pm 0.04$ $0.705\pm 0.04$ $4.91\pm 0.47$ $0.521\pm 0.06$ $0.755\pm 0.03$ $5.51\pm 1.10$ $0.693\pm 0.06$ $0.848\pm 0.07$ U Approximated GED [20] $\underline{9.42}$ 0.116{0.116} 0.176{0.176} 10.44{10.44} 0.285{0.285} 0.407{0.407} $\underline{11.96}$ 0.250{0.250} 0.277{0.277} EUGENE [6] $11.27\pm 0.31$ $0.260\pm 0.03$ $0.333\pm 0.05$ $\underline{9.91\pm 1.55}$ $0.354\pm 0.14$ $0.523\pm 0.12$ $12.27\pm 1.09$ $0.379\pm 0.09$ $0.423\pm 0.08$ GEDAN($\lambda=1$) $\mathbf{6.00\pm 0.10}$ $\mathbf{0.433\pm 0.04}$ $\mathbf{0.658\pm 0.04}$ $\mathbf{8.65\pm 0.94}$ $\mathbf{0.513\pm 0.05}$ $\mathbf{0.741\pm 0.02}$ $\mathbf{5.56\pm 0.55}$ $\mathbf{0.650\pm 0.06}$ $\mathbf{0.839\pm 0.05}$
D.3 Additional Results on Large Graphs
One of the key points raised concerns the possibility of adopting an unsupervised approach to approximate the Graph Edit Distance (GED). As highlighted by Neuhaus et al. (2006a), increasing the graph size inevitably reduces accuracy, since the problem is NP-hard and approximations are required, which in turn lead to suboptimal solutions.
Similarly, in our setting, larger graphs make it increasingly difficult to identify the optimal matching, as the number of candidate pairs grows with the number of nodes. To assess scalability, we consider graphs of up to 32 nodes from the BZR dataset (Morris et al., 2020) using a $64\times 64$ model, graphs of up to 64 nodes from the Mutagenicity and NCI1 datasets with a $128\times 128$ model, and finally the Synthie dataset (average 95 nodes) with a $256\times 256$ model. Despite relying on relatively few graphs due to the training cost, our approach remains consistently faster than EUGENE (Bommakanti et al., 2024).
To estimate GED, we adopt the algorithm proposed by Fischer et al. (2017), implemented in the GMatch4Py library, and also report the GED baseline, a strict lower bound defined as the absolute difference between the number of nodes and edges in the graphs.
Since our model trains more efficiently, we allow it to run for a sufficient number of epochs so that its overall training cost matches the prediction cost of EUGENE (Bommakanti et al., 2024), thereby ensuring a fair comparison. In cases where the number of nodes is particularly large, EUGENE (Bommakanti et al., 2024) failed to provide results within a reasonable time; in such scenarios, we retain the comparison with the method of Fischer et al. (2017).
Figure 12: (a) Comparison with EUGENE and Fischer et al. (2017). The black line indicates the interpolation of Fischer et al. (2017) maxima. (b) Training dynamics of GEDAN, showing the loss, $\tau$, and $\rho$. Training for 20 epochs requires 27 minutes 5 seconds, compared to 26 minutes 21 seconds for EUGENE (Bommakanti et al., 2024).
Figure 13: (a) Comparison with EUGENE and Fischer et al. (2017). The black line indicates the interpolation of Fischer et al. (2017) maxima. (b) Training dynamics of GEDAN, showing the loss, $\tau$, and $\rho$. Training for 20 epochs requires 6 minutes 21 seconds, compared to 6 minutes 44 seconds for EUGENE.
Figure 14: (a) Comparison with Fischer et al. (2017). The black line indicates the interpolation of Fischer et al. (2017) maxima. (b) Training dynamics of GEDAN, showing the loss, $\tau$, and $\rho$. Training for 20 epochs requires 2 minutes 56 seconds.
Figure 15: (a) Comparison with Fischer et al. (2017). The black line indicates the interpolation of Fischer et al. (2017) maxima. (b) Training dynamics of GEDAN, showing the loss, $\tau$, and $\rho$. Training for 20 epochs requires 5 minutes 4 seconds.
Appendix E Details of Edit Cost Optimization